# Gaze3P: Gaze-Based Prediction of User-Perceived Privacy

### v1

@@ Appendix B Cognitive Theories and Privacy
(iv) The heuristic-systematic model (Gharib, 2024) suggest that a heuristic (i.e. quick judgment) may guide the first decision, followed by a deeper systematic analysis to refine privacy preferences.
@@
Appendix C Adverserial Perspective on Differential Privacy
We want to briefly motivate the definition of the adversarial advantage for differential private mechanism we use in Section 2 and Section 6.
@@ Appendix C Adverserial Perspective on Differential Privacy
The adversary wins the security game if $b^{\prime}=b$. The advantage of $\mathcal{A}$ is defined as $\operatorname{adv}_{\mathcal{A}}\coloneqq 2p-1$, where $p$ is the maximal probability that $\mathcal{A}$ wins for any $S$. Note that the adversary in this game is exceptionally strong since it only needs to win the game for one specific output set $S$ (no matter how unlikely $S$ itself is). The setup is nevertheless relevant, since in real-world use cased unlikely outputs might nevertheless occur and even then the privacy of the input data should be preserved.
In slight abuse of notation we also denote by $\operatorname{adv}_{\mathcal{A}}$ the maximal advantage achieved by a ppt. adversary $\mathcal{A}$. We want to determine an upper bound on the advantage. Given the limited information available the most successful adversary $\mathcal{A}$ outputs $b^{\prime}=0$ if $\Pr(0|S)\geq\Pr(1|S)$ and $b^{\prime}=1$ otherwise. Let w.l.o.g. $\Pr(0|S)\geq\Pr(1|S)$. Hence this adversary has advantage $\operatorname{adv}_{\mathcal{A}}=2\Pr(b=0|S)-1$. If the mechanism satisfies $\varepsilon$-differential privacy, Eq. 1 implies $\Pr(S|b=0)\leq e^{\varepsilon}\Pr(S|b=1)$. But $\Pr(S|b=i)=\frac{\Pr(S,b=i)}{\Pr(b=i)}=2\Pr(S,b=i)=2\Pr(S)\Pr(b=i|S)$ for $i=0,1$. Thus $\Pr(S|b=0)\leq e^{\varepsilon}\Pr(S|b=1)\Rightarrow\Pr(b=0|S)\leq e^{\varepsilon}\Pr(b=1|S)=e^{\varepsilon}(1-\Pr(b=0|S))\Rightarrow\Pr(b=0|S)\leq\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. We conclude that $\operatorname{adv}_{\mathcal{A}}=2\Pr(0|S)-1\leq\frac{e^{\varepsilon}}{1+e^{\varepsilon}}-1=\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}$.

### v2

@@ Appendix D Cognitive Theories and Privacy
(iv) The heuristic-systematic model (Gharib, 2024) suggests that a heuristic (i.e., quick judgment) may guide the first decision, followed by a deeper systematic analysis to refine privacy preferences.
@@
Appendix E Adverserial Perspective on Differential Privacy
We want to briefly motivate the definition of the adversarial advantage for the differential private mechanism we use in Section 2 and Section 6.
@@ Appendix E Adverserial Perspective on Differential Privacy
The adversary wins the security game if $b^{\prime}=b$. The advantage of $\mathcal{A}$ is defined as $\operatorname{adv}_{\mathcal{A}}\coloneqq 2p-1$, where $p$ is the maximal probability that $\mathcal{A}$ wins for any $S$. Note that the adversary in this game is exceptionally strong since it only needs to win the game for one specific output set $S$ (no matter how unlikely $S$ itself is). The setup is nevertheless relevant, since in real-world use cases, unlikely outputs might nevertheless occur, and even then, the privacy of the input data should be preserved.
In a slight abuse of notation, we also denote by $\operatorname{adv}_{\mathcal{A}}$ the maximal advantage achieved by a ppt. adversary $\mathcal{A}$. We want to determine an upper bound on the advantage. Given the limited information available the most successful adversary $\mathcal{A}$ outputs $b^{\prime}=0$ if $\Pr(b=0|S)\coloneqq\Pr(D_{0}|M(D_{0})\in S)\geq\Pr(D_{1}|M(D_{1})\in S)\eqqcolon\Pr(b=1|S)$ and $b^{\prime}=1$ otherwise. Let w.l.o.g. $\Pr(b=0|S)\geq\Pr(b=1|S)$. Hence, this adversary has an advantage $\operatorname{adv}_{\mathcal{A}}=2\Pr(b=0|S)-1$. If the mechanism satisfies $\varepsilon$-differential privacy, Eq. 1 implies $\Pr(S|b=0)\leq e^{\varepsilon}\Pr(S|b=1)$. But $\Pr(S|b=i)=\frac{\Pr(S,b=i)}{\Pr(b=i)}=2\Pr(S,b=i)=2\Pr(S)\Pr(b=i|S)$ for $i=0,1$. Thus $\Pr(S|b=0)\leq e^{\varepsilon}\Pr(S|b=1)\Rightarrow\Pr(b=0|S)\leq e^{\varepsilon}\Pr(b=1|S)=e^{\varepsilon}(1-\Pr(b=0|S))\Rightarrow\Pr(b=0|S)\leq\frac{e^{\varepsilon}}{1+e^{\varepsilon}}$. We conclude that $\operatorname{adv}_{\mathcal{A}}=2\Pr(0|S)-1\leq\frac{e^{\varepsilon}}{1+e^{\varepsilon}}-1=\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}$.